# Metric Differential Privacy at the User-Level

### v1

@@ 2.2. Earth-Mover’s Distance
Next we present the Birkhoff-Von Neumann Theorem which is useful in our privacy analysis in Sec. 4.2. The theorem states that if both $P$ and $Q$ are empirical distributions with the same number of points, then the $d_{\textsf{EM}}$ between them is the cost of the coupling that moves the entire mass in each point to the same destination:
@@
[Birkhoff-Von Neumann Theorem (Konig, 2001), Lemma A.1 in (Fernandes et al., 2019)): For two datasets $K=\{x_{1},\ldots,x_{m}\}$ and $K^{\prime}=\{y_{1},\ldots,y_{m}\}$, there is a permutation $\pi:[m]\rightarrow[m]$ such that
(3) $d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})=\frac{1}{m}\sum_{i=1}^{m}d_{\mathcal{X}}(x_{i},y_{\pi(i)}).$
3. Definition of $d_{\textsf{EM}}$-DP
In this section, we introduce our generalization of metric DP to the user-level. We start with the local model. We use the $d_{\textsf{EM}}$ metric to measure the distance between two datasets $K,K^{\prime}$ since it captures the intuition that the changes which move smaller amounts of data by smaller distances are more sensitive (as discussed in Sec. 1).
@@ 3. Definition of $d_{\textsf{EM}}$-DP
(4) $\Pr[\mathcal{M}(K)=O]\leq e^{\alpha d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})}\Pr[\mathcal{M}(K^{\prime})=O]+\delta.$
@@
In the central model, our goal is to protect changes in a single user’s dataset, transitioning from $K_{i}$ to $K^{\prime}_{i}$, with a privacy guarantee that depends on $d_{\textsf{EM}}(\tilde{K},\tilde{K}_{i})$. We consider the bounded data setting where each dataset $K_{i}$ has a publicly known fixed size $m$.
@@
Let $K_{G}=K_{1}\cup\cdots\cup K_{n}$ denote a global dataset from $n$ users where $\forall i\in[n],K_{i}\in\mathcal{X}^{m}$. We say $K_{G}\sim K^{\prime}_{G}$ if $K^{\prime}_{G}$ can be obtained from $K_{G}$ by changing the dataset of a single user $U_{i}$ from $K_{i}$ to $K_{i}^{\prime}$. We say a mechanism $\mathcal{M}$ satisfies $(\alpha,\delta)$-bounded $d_{\textsf{EM}}$-DP if, for all $K_{G},K_{G}^{\prime}$ such that $K_{G}\sim K^{\prime}_{G}$, and all outputs $O$, we have
@@
It is important to note that the above definition cannot be directly translated to the unbounded data setting. This limitation arises from the fact that if each $K_{i}$ is allowed to have an arbitrary size, then changing a single $K_{i}$ could potentially change the entirety of $K_{G}$ in the worst-case (where user $U_{i}$ contributes the entire global dataset). This essentially reduces the central model (Def. 3.2) to the local model (Def. 3.1). We circumvent this challenge and provide a privacy definition for the undounded data setting in Sec. 5, by controlling the amount of data from each user.
Setting the Privacy Parameters. There are some semantic differences between the parameter $\alpha$ in Definitions 3.1 and 3.2, and $\varepsilon$ in Definitions 2.1 and 2.2. The privacy parameter $\varepsilon$ is unitless. On the other hand, $\alpha$ is not unitless – it has a unit inversely proportional to $d_{\textsf{EM}}$. While $\varepsilon\gg 1$ is usually not considered acceptable for standard DP, it is not unreasonable to set $\alpha\gg 1$ in our case. This is acceptable if a strong privacy guarantee is needed only for input pairs that are close to each other since $d_{\textsf{EM}}(\cdot,\cdot)<1$. For all $q,\tau\in[0,1]$, let $\mathcal{E}(q,\tau)$ refer to the minimum privacy parameter that is acceptable over all data changes of the form
@@
Then, $\alpha$ may be set as $\alpha=\inf_{q,\tau\in[0,1]}\tfrac{\mathcal{E}(q,\tau)}{q\tau}$, and we can verify that Definition 3.1 will protect an input pair with the corresponding budget $\mathcal{E}(q,\tau)$. The parameter $\delta$ has the same interpretation as in standard DP, and should be set $\delta\ll\frac{1}{poly(n)}$.
Concrete Example. Throughout this paper, we consider a dataset of $n=10^{5}$ users, each of whom contributes $m=10^{3}$ location data points over the period of a month. We use the length of the shortest path on Earth’s surface as our metric $d_{\mathcal{X}}$. Suppose we want to protect a user’s location over any particular day within a radius of $1000$ miles, and the user’s location over the entire time period within a distance of $100$ miles. In the normalized metric space, these distances are $q_{1}=0.08$ and $q_{2}=0.008$, respectively [^5] The maximum surface distance between two points on Earth is $\approx 12930$ miles.. They correspond to a fraction $\tau_{1}=\frac{1}{30}$ and $\tau_{2}=1$ of the metric space changing, respectively. Suppose we want to protect both of these inputs with privacy parameter $\varepsilon=0.2$. Hence, we set $\alpha=\min\left\{\frac{\varepsilon}{\tau_{1}q_{1}},\frac{\varepsilon}{\tau_{2}q_{2}}\right\}=25.$ This value is much higher than typical privacy parameters used in DP, and yet it is able to adequately protect the desired inputs. Finally, we will set $\delta=10^{-12}$.
4. Mechanisms for $d_{\textsf{EM}}$-DP
Now, we describe our mechanisms for releasing queries under $d_{\textsf{EM}}$-DP. Throughout this section, we focus on the bounded data setting, and consider both the local and central models. In Sec. 4.1, we show how to bound the sensitivity of linear queries, which can then be released with the addition of calibrated noise. Then, in Sec. 4.2, we show that we can release a noisy representation of $\tilde{K}$ under $d_{\textsf{EM}}$-DP by applying any $d_{\mathcal{X}}$-DP mechanism to each item in $K$, and aggregating the outputs.
4.1. Linear Queries
A non-adaptive linear query on a dataset $K$ computes the value of $F\tilde{K}$, where $F\in\mathbb{R}^{d\times|\mathcal{X}|}$ is a matrix with $d$ rows. The linearity comes from the linear transformation $F$; our linear queries are normalized since they operate on $\tilde{K}$ rather than $K$. Such normalized queries can be used for answering the fraction of users satisfying a predicate (Blum et al., 2013). Nevertheless, one can estimate the non-normalized query by multiplying by an estimate of $|K|$.
@@ 4.1. Linear Queries
Naively, it is intractible to compute this sensitivity since there are exponentially many datasets of a given size. Additionally, this sensitivity might not always be bounded. For instance, consider two points $x,x^{\prime}$ that are close in $\mathcal{X}$, but $f(x)$ is very far from $f(x^{\prime})$. In this case, we cannot bound $\Delta_{\mathsf{EM}}$, since the $K$ and $K^{\prime}$ which put all their mass on $x$ and $x^{\prime}$, respectively, will have $\frac{\|q_{f}(K)-q_{f}(K^{\prime})\|}{d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})}=\|f(x)-f(x^{\prime})\|$. However, if $f$ is $\ell$-Lipschitz, meaning
$\max_{x,x^{\prime}\in\mathcal{X}}\frac{\|f(x)-f(x^{\prime})\|}{d_{\mathcal{X}}(x,x^{\prime})}\leq\ell,$
then it is possible to bound $\Delta_{\mathsf{EM}}(q_{f})$ using $\ell$. We do this by observing that, for any coupling between $\tilde{K},\tilde{K}^{\prime}$, each mass that moves a distance $d$ may change $q_{f}$ by up to $\ell d$ (based on Eq. (5)). This allows us to compute the following bound.
@@
Remarks. The above theorem implies that a reasonable upper bound for $\Delta_{\mathsf{EM}}(q_{f})$ can be made when the query function $f$ is smooth in terms of $d_{\mathcal{X}}$. In Sec. 6.1 we outline a special type of linear query for which this is the case. Additionally, the aforementioned example illustrates that this sensitivity analysis is tight. This means that $d_{\mathcal{X}}$, in addition to defining the privacy semantics, also influences the types of queries that can be answered with good utility.
Proof Sketch. Let $C$ be the minimum-cost coupling between $\tilde{K},\tilde{K}^{\prime}$. $\|\mathbb{E}_{x\sim\tilde{K}}[f(x)]-\mathbb{E}_{x\sim\tilde{K}^{\prime}}[f(x)]\|$ can then be bounded by transporting $\tilde{K}^{\prime}$ onto $\tilde{K}$ times the amount that $f$ can change when each mass is transported, which is atmost $\ell d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})$. ∎
Using the upper bound on $\Delta_{\mathsf{EM}}(q_{f})$, we follow a well-known approach for privately releasing a point with known sensitivity under a norm: Sample a point $U$ uniformly from the ball $\{x\in\mathbb{R}^{d}:\|x\|=1\}$, and release $q_{f}+\ell gU$, where $g\sim\Gamma(d,\frac{\omega}{\alpha})$, the Gamma distribution with shape $d$ and scale $\frac{\omega}{\alpha}$ (Hardt and Talwar, 2010). Here, $\omega$ is a scale parameter that may be different in the central or local model, since the sensitivity of $f$ is less in the bounded central model. This mechanism, PrivEMDLinear, is outlined in Alg. 1. Combining Thm. 4.1 with a standard privacy analysis, we can show that PrivEMDLinear satisfies $(\alpha,0)$-$d_{\textsf{EM}}$ DP.
@@
PrivEMDLinear (Alg. 1) with scale $\omega=\frac{1}{\alpha}$ satisfies $(\alpha,0)$-unbounded local $d_{\textsf{EM}}$-DP and with scale $\omega=\frac{1}{\alpha n}$ satisfies $(\alpha,0)$-bounded central $d_{\textsf{EM}}$-DP.
Remarks. When using the $1$-norm, PrivEMDLinear becomes the multidimensional Laplace mechanism. We may instantiate PrivEMDLinear with any noise mechanism that preserves $\alpha\|\cdot\|_{p}$-metric DP in the space $\mathbb{R}^{d}$. In particular, under the $2$-norm, we can add Gaussian noise of width $\frac{\omega\sqrt{1.25\ln(1/\delta)}}{\alpha}$ (Dwork et al., 2014), and this will give better utility than adding noise based on the Gamma distribution at the cost of $(\alpha,\delta)$-bounded local $d_{\textsf{EM}}$ DP.
@@
4.2. Unordered Release of Item-wise Queries
We now consider the problem of directly releasing a private query applied to each item in $K$. This can provide a more fine-grained result than the aforementioned linear queries, which outputs the average over all the items. We release the query results as an unordered list to take advantage of the fact that subsequent computation (such as, aggregation) often does not depend on the ordering of the data (Feldman et al., 2022). Specifically, our second mechanism PrivEMDItemWise applies a mechanism $\mathcal{A}$, which satisfies $\alpha_{0}$-$d_{\mathcal{X}}$ DP, to each item individually. We use $\mathcal{A}$ as a black-box making PrivEMDItemWise completely general-purpose. For example, one could let $\mathcal{A}$ be a private item-release mechanism [^6] A number of metric DP mechanisms for releasing items in specific applications are mentioned in Sec. 7. and use PrivEMDItemWise to form a histogram of the dataset. $\mathcal{A}$ could also be a classifer, and PrivEMDItemWise can then release a simplified representation of the dataset.
Once PrivEMDItemWise applies $\mathcal{A}$ to each element in the dataset, it shuffles the results (to remove any ordering of the data) and outputs the shuffled list. This appears in Alg. 2, and a precursor to this algorithm appeared in (Fernandes et al., 2019).

### v2

@@ 2.2 Earth-Mover’s Distance
Next, we present the Birkhoff-Von Neumann theorem which is useful in our privacy analysis in Section 4.2. The theorem states that if both $P$ and $Q$ are empirical distributions with the same number of points, then the $d_{\textsf{EM}}$ between them is the cost of the coupling that moves the entire mass in each point to the same destination:
@@
[Birkhoff-Von Neumann Theorem Konig (2001), Lemma A.1 in Fernandes et al. (2019)): For two datasets $K=\{x_{1},\ldots,x_{m}\}$ and $K^{\prime}=\{y_{1},\ldots,y_{m}\}$, there is a permutation $\pi:[m]\rightarrow[m]$ such that
$d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})=\frac{1}{m}\sum_{i=1}^{m}d_{\mathcal{X}}(x_{i},y_{\pi(i)}).$ (3)
3 Definition of $d_{\textsf{EM}}$-DP
In this section, we introduce our generalization of metric DP to the user-level. We start with the local model. We use the $d_{\textsf{EM}}$ metric to measure the distance between two datasets $K,K^{\prime}$ since it captures the intuition that the changes which move smaller amounts of data by smaller distances are more sensitive (as discussed in Section 1).
@@ 3 Definition of $d_{\textsf{EM}}$-DP
$\Pr[\mathcal{M}(K)=O]\leq e^{\alpha d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})}\Pr[\mathcal{M}(K^{\prime})=O]+\delta.$ (4)
@@
In the central model, our goal is to protect changes in a single user’s dataset, transitioning from $K_{i}$ to $K^{\prime}_{i}$, with a privacy guarantee that depends on $d_{\textsf{EM}}(\tilde{K}_{i},\tilde{K}_{i}^{\prime})$. We consider the bounded data setting where each dataset $K_{i}$ has a publicly known fixed size $m$.
@@
Let $K_{G}=K_{1}\cup\cdots\cup K_{n}$ denote a global dataset from $n$ users where $\forall i\in[n],K_{i}\in\mathcal{X}^{m}$. We say $K_{G}\sim K^{\prime}_{G}$ if $K^{\prime}_{G}$ can be obtained from $K_{G}$ by changing the dataset at a single index $i$ from $K_{i}$ to $K_{i}^{\prime}$. We say a mechanism $\mathcal{M}$ satisfies $(\alpha,\delta)$-bounded central $d_{\textsf{EM}}$-DP if, for all $K_{G},K_{G}^{\prime}$ such that $K_{G}\sim K^{\prime}_{G}$, and all outputs $O$, we have
@@
It is important to note that the above definition cannot be directly translated to the unbounded data setting. This limitation arises from the fact that if each $K_{i}$ is allowed to have an arbitrary size, then changing a single $K_{i}$ could potentially change the entirety of $K_{G}$ in the worst-case (where user $U_{i}$ contributes the entire global dataset). This essentially reduces the central model (Def. 3.2) to the local model (Def. 3.1). We circumvent this challenge and provide a privacy definition for the unbounded data setting in Section 5, by controlling the amount of data from each user.
Setting the Privacy Parameters. There are some semantic differences between the parameter $\alpha$ in Defns. 3.1 and 3.2, and $\varepsilon$ in Defns. 2.1 and 2.2. The privacy parameter $\varepsilon$ is unitless. On the other hand, $\alpha$ is not unitless – it has a unit inversely proportional to $d_{\textsf{EM}}$. While $\varepsilon\gg 1$ is usually not considered acceptable for standard DP, it is not unreasonable to set $\alpha\gg 1$ in our case. This is acceptable if a strong privacy guarantee is needed only for input pairs that are close to each other since $d_{\textsf{EM}}(\cdot,\cdot)<1$. For all $q,\tau\in[0,1]$, let $\mathcal{E}(q,\tau)$ refer to the minimum privacy parameter that is acceptable over all data changes of the form
@@
Then, $\alpha$ may be set as $\alpha=\inf_{q,\tau\in[0,1]}\tfrac{\mathcal{E}(q,\tau)}{q\tau}$, and we can verify that Defn. 3.1 will protect an input pair with the corresponding budget $\mathcal{E}(q,\tau)$. The parameter $\delta$ has the same interpretation as in standard DP, and should be set $\delta\ll\frac{1}{poly(n)}$.
Concrete Example. Throughout this paper, we consider a dataset of $n=10^{5}$ users, each of whom contributes $m=10^{3}$ location data points over the period of a month. We use the length of the shortest path on earth’s surface as our metric $d_{\mathcal{X}}$. Suppose we want to protect a user’s location over any particular day within a radius of $1000$ miles, and the user’s location over the entire time period within a distance of $100$ miles. In the normalized metric space, these distances are $q_{1}=0.08$ and $q_{2}=0.008$, respectively [^3] The maximum surface distance between two points on Earth is $\approx 12930$ miles.. They correspond to a fraction $\tau_{1}=\frac{1}{30}$ and $\tau_{2}=1$ of the metric space changing, respectively. Suppose we want to protect both of these inputs with privacy parameter $\varepsilon=0.2$. Hence, we set $\alpha=\min\left\{\frac{\varepsilon}{\tau_{1}q_{1}},\frac{\varepsilon}{\tau_{2}q_{2}}\right\}=25.$ This value is much higher than typical privacy parameters used in DP, and yet it is able to adequately protect the desired inputs. Finally, we will set $\delta=10^{-12}$ in our examples.
4 Mechanisms for $d_{\textsf{EM}}$-DP
Now, we describe our mechanisms for releasing queries under $d_{\textsf{EM}}$-DP. Throughout this section, we focus on the bounded data setting, and consider both the local and central models by considering a general dataset $K$ which will be set to be $K_{i}$ in the local model and $K_{G}$ in the central one. In Section 4.1, we show how to bound the sensitivity of linear queries, which can then be released with the addition of calibrated noise. Then, in Section 4.2, we show that we can release a noisy representation of $\tilde{K}$ under $d_{\textsf{EM}}$-DP by applying any $d_{\mathcal{X}}$-DP mechanism to each item in $K$, and shuffling the outputs. Full proofs for this section appear in Appendix B.
4.1 Linear Queries
A non-adaptive linear query on a dataset $K$ computes the value of $F\tilde{K}$, where $F\in\mathbb{R}^{d\times|\mathcal{X}|}$ is a matrix with $d$ rows. The linearity comes from the linear transformation $F$; our linear queries are normalized since they operate on $\tilde{K}$ rather than $K$. Such normalized queries can be used for answering the fraction of users satisfying a predicate Blum et al. (2013). Nevertheless, one can estimate the non-normalized query by multiplying by an estimate of $|K|$.
@@ 4.1 Linear Queries
Naively, it is intractible to compute this sensitivity since there are exponentially many datasets of a given size. Additionally, this sensitivity might not always be bounded. For instance, consider two points $x,x^{\prime}$ that are close in $\mathcal{X}$, but $f(x)$ is very far from $f(x^{\prime})$. In this case, we cannot bound $\Delta_{\mathsf{EM}}$, since the $K$ and $K^{\prime}$ which put all their mass on $x$ and $x^{\prime}$, respectively, will have $\frac{\|q_{f}(K)-q_{f}(K^{\prime})\|}{d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime})}=\|f(x)-f(x^{\prime})\|$. We may exclude this case by assuming that $f$ is $\ell$-Lipschitz, meaning
$\max_{x,x^{\prime}\in\mathcal{X}}\frac{\|f(x)-f(x^{\prime})\|}{d_{\mathcal{X}}(x,x^{\prime})}\leq\ell.$
It turns out that Lipschitzness is precisely the property needed in order to bound $\|q_{f}(\tilde{K})-q_{f}(\tilde{K}^{\prime})\|$ while effectively accommodating the underlying coupling between $\tilde{K},\tilde{K}^{\prime}$. This is a novel aspect of our analysis that has not been explored by previous sensitivity analysis.
@@
Remarks. The above result stands in contrast with traditional sensitivity arguments, which typically assume the weaker condition that $f$ is merely bounded by $\ell$. Although our Lipschitz assumption is stronger, it is satisfied by many queries of interest, such as
  • Average. If $\mathcal{X}\subseteq\mathbb{R}^{d}$, then taking $f$ to be the identity function will cause $q_{f}$ to simply be the average of the elements in $K$.
  • Kernel Smoothed Region Queries. If $R\subseteq\mathcal{X}$ is a region of interest, then $f(x)=\textbf{1}[x\in R]$ will yield a $q_{f}$ computing the fraction of elements that lie in $R$. Using a smooth interpolation of $\textbf{1}[x\in R]$, such as kernel smoothers Hastie et al. (2009), will give a smooth approximation to this value.
  • Similarity Queries. If $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $c\in\mathbb{R}^{d}$ is a query vector, then letting $f(x)=\langle x,c\rangle$ will return the average similarity of each vector in $K$ with $c$, with similarity measured by the dot product. This is an instance of a linear embedding query which we will study in detail in Section 6.1.
Additionally, the aforementioned example illustrates that this sensitivity analysis is tight. This means that $d_{\mathcal{X}}$, in addition to defining the privacy semantics, also influences the types of queries that can be answered with good utility.
Proof Sketch. At a high level, consider moving $K$ onto $K^{\prime}$ one point at a time. The Lipschitz assumption allows us to bound the change in $q_{f}(K)$ in terms of how far the point is moving, which translates to the $d_{\textsf{EM}}$ distance. More formally, let $C$ be a minimum cost coupling between $\tilde{K},\tilde{K}^{\prime}$. We may write
$\|q_{f}(\tilde{K})-q_{f}(\tilde{K}^{\prime})\|=\left\|\mathbb{E}_{x\sim\tilde{K}}[f(x)]-\mathbb{E}_{y\sim\tilde{K}^{\prime}}[f(y)]\right\|.$
We can view $x$ and $y$ as jointly generated by $C$, and the above expression becomes $\|\mathbb{E}_{x,y\sim C}[f(x)-f(y)]\|$. By the triangle inequality, we know this is at most
$\|q_{f}(\tilde{K})-q_{f}(\tilde{K}^{\prime})\|\leq\mathbb{E}_{x,y\sim C}[\|f(x)-f(y)\|],$
and by the Lipschitz assumption, we may upper bound this by
$\ell\cdot\mathbb{E}_{x,y\sim C}[d_{\mathcal{X}}(x,y)]=\ell\times d_{\textsf{EM}}(\tilde{K},\tilde{K}^{\prime}),$
which completes the proof. ∎
Using the upper bound on $\Delta_{\mathsf{EM}}(q_{f})$, we follow a well-known approach for privately releasing a point with known sensitivity under a norm: sample a point $U$ uniformly from the ball $\{x\in\mathbb{R}^{d}:\|x\|=1\}$, and release $q_{f}+\ell gU$, where $g\sim\Gamma(d,\frac{\omega}{\alpha})$ is the Gamma distribution with shape $d$ and scale $\frac{\omega}{\alpha}$ (Hardt and Talwar, 2010). Here, $\omega$ is a scale parameter that may be different in the central or local model, since the sensitivity of $f$ is less in the bounded central model. This mechanism, PrivEMDLinear, is outlined in Algorithm 1. Combining Theorem 4.1 with a standard privacy analysis, we can show that PrivEMDLinear satisfies $(\alpha,0)$-$d_{\textsf{EM}}$ DP.
@@
PrivEMDLinear (Algorithm 1) with scale $\omega=\frac{1}{\alpha}$ satisfies $(\alpha,0)$-unbounded local $d_{\textsf{EM}}$-DP and with scale $\omega=\frac{1}{\alpha n}$ satisfies $(\alpha,0)$-bounded central $d_{\textsf{EM}}$-DP.
Remarks. When using the $1$-norm, PrivEMDLinear becomes the multidimensional Laplace mechanism. We may instantiate PrivEMDLinear with any noise mechanism that preserves $(\alpha,\delta)$-local $\|\cdot\|_{p}$-DP in the space $\mathbb{R}^{d}$. In Section 6.1, we will instantiate PrivEMDLinear using Gaussian noise of width $\frac{\omega\sqrt{1.25\ln(1/\delta)}}{\alpha}$ Dwork et al. (2014), which will give the proper error dependence on $d$ under the $2$-norm.
@@
4.2 Unordered Release of Item-wise Queries
We now consider the problem of directly releasing a private query applied to each item in $K$. This can provide a more fine-grained result than the aforementioned linear queries, which outputs the average over all the items. We release the query results as an unordered list to take advantage of the fact that subsequent computation (such as, aggregation) often does not depend on the ordering of the data Feldman et al. (2022). Specifically, our second mechanism PrivEMDItemWise applies a mechanism $\mathcal{A}$, which satisfies $(\alpha_{0},0)$-$d_{\mathcal{X}}$-DP, to each item individually. We use $\mathcal{A}$ as a black-box making PrivEMDItemWise completely general-purpose. For example, one could let $\mathcal{A}$ be a private item-release mechanism (see Section 7 for some examples) and use PrivEMDItemWise to form a histogram of the dataset. $\mathcal{A}$ could also be a classifer, and PrivEMDItemWise can then release a simplified representation of the dataset. Once PrivEMDItemWise applies $\mathcal{A}$ to each element in the dataset, it shuffles the results (to remove any ordering of the data) and outputs the shuffled list. This appears in Algorithm 2, and a precursor appeared in Fernandes et al. (2019).
As PrivEMDItemWise does not hide the size of $K$, we show it satisfies bounded $d_{\textsf{EM}}$-DP. We use the following argument: for a neighboring dataset $K^{\prime}=\{x_{1}^{\prime},\ldots,x_{m}^{\prime}\}$, by Lemma 2.1 there exists a permutation $\pi:[m]\rightarrow[m]$ satisfying Eq. (3). Observe that we release the query responses in an unordered fashion by explicitly shuffling them. This allows us to pair up the element $x_{i}$ with $x_{\pi(i)}$ and analyze the privacy guarantee of releasing $\mathcal{A}(x_{1}),\ldots,\mathcal{A}(x_{m})$ versus $\mathcal{A}(x_{\pi(i)}^{\prime}),\cdots,\mathcal{A}(x_{\pi(m)}^{\prime})$. Prior work does this with composition Fernandes et al. (2019).